OpenStudy (anonymous):

If f(x) = 4x – 5 and g(x) = 3 – 2x, find f(x) + g(x).

OpenStudy (haseeb96):

just add it like simple addition and put all these values together f(x)+g(x)=4x-5 +3-2x =2x-2 = 2(x-1) is the correct answer

OpenStudy (anonymous):

So you group them first then add? And im confused on the answer is it 2x-2? or 2(x-1)?

OpenStudy (anonymous):

A. 2x – 2 B. 6x – 2 C. 2x – 8 D. 6x – 8

OpenStudy (haseeb96):

A is the correct answer 2(x-1) is the simplest form of 2x-2

OpenStudy (anonymous):

Thank you.

OpenStudy (haseeb96):

welcome